OpenStudy (anonymous):

find the derivative of f(x)=-9/x at x=-8

zepdrix (zepdrix):

Have you learned your shortcuts for taking derivatives yet? Or should we use the `Limit Definition for the Derivative`?

OpenStudy (anonymous):

I haven't learned the shortcuts yet

OpenStudy (anonymous):

isn't it 9/64 ?

zepdrix (zepdrix):

Using the limit definition: \[\large f'(-8)=\frac{\dfrac{-9}{-8+h}-\dfrac{-9}{-8}}{h}\] Which simplifies toooo, yah 9/64 sounds right, good job.

OpenStudy (anonymous):

thank you!(: